OpenStudy (anonymous):

MEDAL AND FAN!!! Given f(x)=x-5/3 solve for f-1(3). HELP PLEASE!

OpenStudy (anonymous):

f-1 is for inverse. You need to find the inverse of the function first then solve.

OpenStudy (anonymous):

Yes i know, but I don't remember how to! can you walk me through it please?

OpenStudy (anonymous):

To do inverse of f(x) = x-5/3..we first make f(x) =y so it will be easy to follow. So our equation is y = x-5/3..do you see this?

OpenStudy (anonymous):

Yes, then you solve for x right?

OpenStudy (anonymous):

No to get the inverse of a function you switch the x and y and solve for y. So in this equation we swap the x and y we end up with x=y-5/3. Do you understand?

OpenStudy (anonymous):

Oh okay. So after you solve for y you make it like equal to f(-1)?

OpenStudy (anonymous):

What will be the inverse function you end up with?

OpenStudy (anonymous):

f-1(x)=3x+5?

OpenStudy (anonymous):

No, If x = y-5/3. What will y= ???

OpenStudy (anonymous):

Uh I don't know. I'm confused :( I'm sorry

OpenStudy (anonymous):

Can you please explain it further?

OpenStudy (anonymous):

The original function is f(x)= x -5/3. We make f(x) =y, and we have y= x - 5/3. We swap x and y and we have x= y -5/3. Now we solve for y by adding 5/3 to both sides giving as y = x+5/3. (THIS IS OUR INVERSE FUNCTION). The question ask us to solve f-1(3) which is the same as f-1(x) which means we plug 3 into the value of x and solve for y and you have your answer.

OpenStudy (anonymous):

OMG THANK YOU! so we plug in that into x=y-5/3 or y-x+5/3?

OpenStudy (anonymous):

Yu plug it into the inverse function

OpenStudy (anonymous):

my answer choices are -1 14 18 21 but my answer was none of those?

OpenStudy (anonymous):

@nelsonjedi???

OpenStudy (anonymous):

Have no idea

OpenStudy (anonymous):

YOU NEED TO WRITE THE EQUATION BETTER.....The original function is f(x)= (x -5)/3. We make f(x) =y, and we have y= (x - 5)/3. We swap x and y and we have x= (y -5)/3. Now we solve for y by multiplying 3 to both sides giving as 3x=y-5. Now add 5 to both sides we get y=3x+5 (THIS IS OUR INVERSE FUNCTION). The question ask us to solve f-1(3) which is the same as f-1(x) which means we plug 3 into the value of x and solve for y and you have your answer.